### v2

# Evaluating Disentangled Representations for Controllable Music Generation

Abstract
Recent approaches in music generation rely on disentangled representations, often labeled as structure and timbre or local and global, to enable controllable synthesis. Yet the underlying properties of these embeddings remain underexplored. In this work, we evaluate such disentangled representations in a set of music audio models for controllable generation using a probing-based framework that goes beyond standard downstream tasks. The selected models reflect diverse unsupervised disentanglement strategies, including inductive biases, data augmentations, adversarial objectives, and staged training procedures. We further isolate specific strategies to analyze their effect. Our analysis spans four key axes: informativeness, equivariance, invariance, and disentanglement, which are assessed across datasets, tasks, and controlled transformations. Our findings reveal inconsistencies between intended and actual semantics of the embeddings, suggesting that current strategies fall short of producing truly disentangled representations, and prompting a re-examination of how controllability is approached in music generation.
Index Terms—  Disentangled representations, controllable music generation, evaluation framework
1 Introduction
Over the past few years, text-to-music generation has achieved high-fidelity audio synthesis [1, 7, 10], with research efforts shifting towards improving controllability [25, 22, 3]. Although text conditioning provides creative flexibility, it does not enforce temporal order or structure [1], and fails to capture lower-level attributes required for precise musical intent [25]. On the other hand, controllability is understood as the ability of a generative model to selectively and predictably manipulate specific attributes of the output (e.g., rhythm or scale) [18]. This directly motivates disentangled representations, commonly described as those in which each component of learned features refers to a semantically meaningful concept [12].
In this context, several methods aim to disentangle music audio into a global embedding, often associated with timbre, and a time-varying embedding, typically linked to structural elements like rhythm or pitch. This separation enables controllable generation and one-shot style transfer [6, 15, 8]. Building on earlier work in representation learning that separates pitch and timbre in isolated musical notes [13, 14], recent approaches propose to disentangle these concepts directly from music audio in an unsupervised manner. This is achieved through architectural constraints [6, 15], timbre-preserving data augmentations [6, 8], adversarial objectives [8], or two-stage training strategies [15, 8]. However, the evaluation of these models tends to focus on the generated output rather than on the learned representations themselves, leaving their suitability for controllable generation uncertain. This raises important questions: What do these representations really capture? Are musical concepts truly disentangled?
Meanwhile, research on foundation models has primarily relied on representation probing as the standard evaluation method for analyzing the information encoded in model embeddings [20]. In the music domain, such frameworks have been applied to evaluate learned representations across several tasks, including key detection, beat tracking, chord recognition, and music tagging [24, 2], as well as to examine how music-theoretic concepts are embedded in the internal representations of generative models [23]. Building on this, Plachouras et al. [20] introduced the synesis framework, originally applied to image and speech models, that assesses representations across multiple axes: informativeness, equivariance, invariance, and disentanglement. Unlike traditional downstream probing, which primarily tests for the presence of task-relevant information in the latent space, this framework examines the structural properties of representations, offering a more holistic view of their quality.
In this work, we adapt this framework to the evaluation of disentangled representations in music audio. Specifically, we: (i) tailor the evaluation protocol to disentangled representations, (ii) apply it to music audio by comparing three unsupervised structure–timbre disentanglement models, and (iii) investigate the effect of specific disentanglement strategies, such as data augmentations and adversarial objectives, on the learned representations.
Our findings reveal high mutual information between the two embeddings, raising questions about the true semantics of these representations and their suitability for tasks such as timbre transfer. The results also prompt a broader reflection on the definition of controllability in music generation, suggesting that current unsupervised strategies fall short of producing representations that consistently support it.
Model Timbre Dims Struct. Dims Temp. Res Data Aug. Arch. Constr. TS Train. Adv. Loss
$S$-VQ-VAE 1024 1024 9 S+P
TS-DSAE 16 16 63 None
AFTER 6 12 86 S+P+T
AFTER-no-aug 6 12 86 S
AFTER-no-adv 6 12 86 S+P+T
Table 1: Configurations of evaluated models, including embedding sizes, resolution, and disentanglement strategies. Data augmentations on the timbre encoder input include segment pairs (S), pitch shifts (P), and time stretches (T).
2 Method
2.1 Models and Disentanglement Strategies
We assess the following models for unsupervised disentanglement of music audio into structure and timbre representations:
  • SS-VQ-VAE [6], designed for style transfer, disentangles content and style through a discrete codebook for content and a style encoder regularized with data augmentations.
  • TS-DSAE [15] extends the Disentangled Sequential Autoencoder [12] with a two-stage training framework that promotes the separation of local and global factors.
  • AFTER [8] aims to disentangle structure and timbre by combining a two-stage training procedure with an adversarial objective and timbre-preserving data augmentations.
Despite their different terminologies, all models are designed or evaluated under the assumption that global embeddings capture timbre or instrument identity, while associating the time-varying embedding with structural elements such as pitch, supporting one-shot timbre transfer. We therefore adopt a structure-timbre terminology throughout this work. We further use AFTER as a testbed to isolate specific disentanglement strategies, namely data augmentations and adversarial loss, in order to study their effect across evaluation axes.
2.2 Probing Methods
Probes are classifiers trained to predict a specific property from intermediate representations of a frozen model [19]. In recent years, they have become a standard tool for analyzing generative model representations in Music Information Retrieval (MIR) [19, 4, 23]. In synesis [20], this corresponds to the axis of informativeness. Formally, given an embedding $z=f(x)$ and a target label $y$, probe $g$ is trained to minimize supervised loss $\mathcal{L}(g(z),y)$, the performance serving as a proxy for the amount of information about $y$ encoded in $z$.
In our setting of disentangled representations, each embedding is probed according to its intended semantics: timbre embeddings on timbre-related tasks such as instrument classification, and structure embeddings on music-theoretic tasks such as multi-pitch estimation and tempo prediction.
Timbre Structure
S. Instr. MPE S. Chords S. Notes S. Tempos
Model Acc ↑ F1 ↑ Acc ↑ Acc ↑ MSE ↓
$S$-VQ-VAE 0.982 0.258 0.462 0.401 0.496
TS-DSAE 0.286 0.133 0.243 0.354 0.187
AFTER 0.284 0.162 0.263 0.311 0.745
AFTER-no-aug 0.260 0.164* 0.266* 0.309 0.716*
AFTER-no-adv 0.266 0.168* 0.251 0.280 0.794
Table 2: Informativeness of timbre and structure embeddings across datasets and downstream probing tasks.
P-Equivariance R-Equivariance
Timbre Structure Timbre Structure
Instr. Change Pitch Shift Time Stretch Instr. Change Pitch Shift Time Stretch
Model MSE ↓ MSE ↓ MSE ↓ Cos sim ↑ Cos sim ↑ Cos sim ↑
$S$-VQ-VAE 0.029 0.127 0.032 0.710 0.823 0.850
TS-DSAE 0.026 0.090 0.079 0.838 0.965 0.974
AFTER 0.028 0.078 0.080 0.770 0.825 0.940
AFTER-no-aug 0.028 0.083 0.076* 0.681 0.892* 0.934
AFTER-no-adv 0.037 0.074* 0.078* 0.653 0.806 0.932
Table 3: P-Equivariance and R-Equivariance evaluation of timbre and structure embeddings across defined transformations.
2.3 Beyond Downstream Tasks
To move beyond downstream prediction, we adopt [20], which assesses representations in terms of their structural response to controlled transformations applied to the input.
Equivariance captures the consistency between transformations in input space and their correspondence in embedding space. Formally, given a transformation $t$ acting on the input $x$, equivariance holds if $z(t(x))\approx\rho_{t}(z(x))$, where $z=f(x)$ is the embedding and $\rho_{t}$ the corresponding action in representation space. We measure this either by predicting transformation parameters (P-equivariance) or by predicting the transformed embedding (R-equivariance). Invariance, in contrast, assesses the stability of embeddings under perturbations that should not alter semantic content, measured by the similarity between embeddings of clean and perturbed inputs.
For disentangled representations, transformations are applied selectively: e.g., structure embeddings are evaluated for equivariance to pitch shifts and invariance to instrument changes, while timbre embeddings are evaluated conversely.
2.4 Measuring Disentanglement
Disentanglement is often measured using metrics such as the Mutual Information Gap (MIG) [5] or DCI [9], which quantify the independence of factors of variation across latent dimensions within a single representation. Alternatively, Plachouras et al. [20] proposed to assess disentanglement by measuring probe performance drops when predicting one factor under controlled transformations of another.
In our setting, where timbre and structure are encoded separately, we extend this principle by comparing probe performance when trained on a single embedding versus on the concatenation of both. Formally, for a task associated with embedding $z_{t}$, we compute:
$\Delta=\big|\text{Acc}(g^{\prime}(z_{t}\oplus z_{s}))-\text{Acc}(g(z_{t}))\big|,$
where Acc denotes probe accuracy, $g$ is the corresponding informativeness probe trained on $z_{t}$, and $g^{\prime}$ is a new probe trained on the concatenation ($\oplus$) of $z_{t}$ and $z_{s}$. A large $\Delta$ indicates that the complementary embedding carries residual information about the task, suggesting entanglement, while values closer to zero imply better embedding separation.
Timbre Structure
Pitch Shift Time Stretch Instr. Change
Model Cos sim ↑ Cos sim ↑ Cos sim ↑
$S$-VQ-VAE 0.667 0.963 0.919
TS-DSAE 0.491 0.993 0.960
AFTER 0.546 0.996 0.960
AFTER-no-aug 0.576* 0.997* 0.955
AFTER-no-adv 0.366 0.984 0.925
Table 4: Invariance of timbre and structure embeddings under different perturbations, reported as cosine similarity.
3 Experimental Setup
3.1 Model Training and Configurations
For a fair comparison between the different models and disentanglement strategies, we retrain all variants using the Slakh2100 [17] dataset, a 145-hour collection of synthetic mixtures with corresponding individual stems. As in [8], mixes and drums are excluded, while the remaining stems are divided between 90% for training and 10% for validation.
The three models SS-VQ-VAE, TS-DSAE, and AFTER are trained using their official repositories [^1] For AFTER, we rely on the real-time model implementation. with default configurations, assumed to be optimal for each disentanglement strategy despite differing embedding sizes. In addition, we train two variants of AFTER to isolate specific disentanglement strategies: one without pitch and tempo augmentations applied to the timbre encoder input (AFTER-no-aug), and one without the adversarial loss (AFTER-no-adv). Table 1 summarizes the different model configurations under study.
3.2 Evaluation Framework
We build on the synesis evaluation framework [20], adapting it [^2] Additional implementation details and an overview figure are available at: https://github.com/lauraibnz/synesis to our setting through the following components.
Probing datasets. We primarily rely on the SynTheory (S.) dataset [19], which provides controlled synthesized variations defined over Western music-theoretic concepts, expected to be present in Slakh2100 in a more complex and naturally entangled form. For timbre, we use SynTheory Instruments, derived from the Chords subset by performing instrument classification (92 MIDI program numbers). For structure, we use SynTheory Notes for root-note classification (12 pitch classes), SynTheory Tempos for tempo regression (161 bpm values), and SynTheory Chords for chord classification (4 chord types). We additionally evaluate the structure embedding on multi-pitch estimation (MPE) to measure its performance on note-level tasks. For this, we train on 4-second excerpts from 36 hours of MAESTRO data [11], and evaluate on the MusicNet test set with corrected alignments [16].
Factors of variation. We consider pitch, tempo, and instrument timbre as factors of variation, and probe equivariance and invariance using three corresponding transformations. Instrument Change, derived from SynTheory, uses identical structures rendered with different instruments and is quantified using the timbre dissimilarity metric of [6], which measures perceptual distance between instrument timbres. This tests timbre equivariance and structure invariance. Pitch Shift and Time Stretch, implemented in synesis, parameterize pitch and tempo variations and are used to evaluate structure equivariance and timbre invariance.
Training details. For standard classification and regression tasks, we follow synesis and train shallow linear probes (SLPs) with early stopping on the frozen embeddings across all evaluation axes. For global tasks on the structure embedding, we apply average pooling over the time dimension. For multi-pitch estimation, we use a two-layer MLP with 512 hidden units and apply a sigmoid activation at the output layer.
Evaluation metrics. Across all evaluation axes, we adopt the default metrics from synesis: accuracy for classification, mean squared error (MSE) for regression, and cosine similarity for embedding similarity. For multi-pitch estimation, we use the mir_eval library [21] to compute F1 scores. The decision threshold is tuned on the validation set and then applied to the test set, with final results averaged at the track level.
Timbre Structure
S. Instr. MPE S. Chords S. Notes S. Tempos
[…]
$S$-VQ-VAE 0.002 0.031 0.311 0.318 0.478
TS-DSAE 0.015 0.016 0.066 0.034 0.174
AFTER 0.068 0.005 0.001 0.009 0.382
AFTER-no-aug 0.097 0.003* 0.048 0.004* 0.458
AFTER-no-adv 0.151 0.056 0.067 0.015 0.298*
Table 5: Disentanglement evaluation across downstream tasks. $\Delta$ measures difference from concatenated embeddings.
4 Results
In all tables, best results are in bold, and asterisk (*) indicates an AFTER variant outperforming the default configuration.
Informativeness. As shown in Table 2, SS-VQ-VAE performs best on instrument classification for timbre, likely due to its larger embedding size. This aligns with [20], which links informativeness to embedding capacity, and the same trend extends to most structure tasks. Remarkably, it also achieves the best multi-pitch estimation performance despite its lower temporal resolution; however, overall performance on this note-level task remains low. In contrast, TS-DSAE is most effective for tempo regression, suggesting that its architecture and training strategy compensate for its smaller embedding size. AFTER remains less consistent, with notably weaker tempo prediction and only minor variant-specific effects.
Equivariance. Table 3 summarizes the equivariance results, revealing a negative correlation with informativeness and embedding size: SS-VQ-VAE, while highly informative, is less equivariant. This trade-off is consistent with observations in the original framework [20]. In contrast, TS-DSAE, despite its compact structure embedding, achieves the strongest R-equivariance across all transformations. AFTER attains intermediate performance, with variants showing reduced timbre equivariance and mixed outcomes for structure.
Invariance. Table 4 presents the invariance results. Timbre embeddings are robust to tempo perturbations but sensitive to pitch shifts. SS-VQ-VAE shows the strongest pitch invariance, followed by AFTER. Its variant without adversarial loss is most vulnerable, while removing pitch and tempo augmentations surprisingly results in higher invariance to both factors. Structure embeddings are broadly invariant to instrument changes, though AFTER again weakens when adversarial loss is removed. Consistent with the original framework [20], invariance appears driven less by capacity and more by training objectives and disentanglement strategies.
Disentanglement. Table 5 summarizes the disentanglement results. SS-VQ-VAE shows little timbre leakage into the structure embedding, but its timbre embedding carries substantial structural information, likely due to its higher capacity and despite data augmentations. AFTER, in contrast, reveals the opposite trend: its structure embedding contains notable timbre information, while the timbre embedding generally carries little pitch information. Additional structural leakage is observed in variants without disentanglement strategies. Across models, timbre embeddings systematically encode tempo, as concatenation consistently improves tempo regression performance. Overall, TS-DSAE appears the most balanced, with concatenation having only marginal effects.
5 Conclusion
We presented a systematic evaluation of disentangled music representations, adapting a recent framework [20] that goes beyond standard downstream probing. Our analysis assessed informativeness, equivariance, invariance, and disentanglement across modern unsupervised architectures for structure–timbre disentanglement in music audio, while also investigating the role of specific disentanglement strategies.
Results show systematic inconsistencies between the intended and actual semantics of the embeddings. We observe asymmetric leakage between timbre and structure, as well as persistent tempo information encoded in timbre embeddings. This limits controllability and reduces suitability for timbre transfer tasks. Moreover, while these models achieve high reconstruction quality, note-level tasks with simple probes fail to extract sufficient information from the embeddings, suggesting such information is encoded in a highly complex way.
More broadly, our evaluation confirms trends observed in the original framework: a trade-off between informativeness and equivariance, with larger embeddings encoding more information but being less transformation-consistent. At the same time, disentanglement strategies such as adversarial losses or data augmentations have the strongest impact on invariance and disentanglement, influencing how well factors are separated rather than how much information is captured.
Finally, while our focus has been on representation-level analysis, controllability also depends on decoding strategies. As noted in [18], disentangled representations are necessary but not sufficient: predictable control requires decoders to preserve factor separation at the signal level. Extending representation-level diagnostics with output-level evaluations, including controlled generation experiments and perceptual listening-based assessments of controllability, remains an important direction for future work.
6 Acknowledgments
This work has been supported by IA y Música: Cátedra en Inteligencia Artificial y Música (TSI-100929-2023-1), funded by the Secretaría de Estado de Digitalización e Inteligencia Artificial and the European Union-Next Generation EU, and IMPA: Multimodal AI for Audio Processing (PID2023-152250OB-I00), funded by the Ministry of Science, Innovation and Universities of the Spanish Government, the Agencia Estatal de Investigación (AEI) and co-financed by the European Union.